OpenStudy (jhannybean):

Determine the point on the plane with equation x+2y+3z=6 that is closest to (1,2,3)

OpenStudy (jhannybean):

i know the short way of finding the distance, which is \[\large D = \frac{\left|ax+by+cd+d\right|}{\sqrt{a^2+b^2+c^2}}\] but my professor wanted us to use the partial derivative method... and the formula for the distance \(\large D(x,y) =f_{xx}f_{yy} -(f_{xy})^2\)

OpenStudy (psymon):

I would think that theres so much of an easier way :/ I would have an idea how to do it if it werent partial deriv method, haha x_X

OpenStudy (usukidoll):

partial derivative in terms of x twice multiplied by partial derivative in terms of y twice

OpenStudy (jhannybean):

Made an error, supposed to say \[\large D= \frac{\left|ax+by+cz+d\right|}{\sqrt{a^2+b^+c^2}}\]

OpenStudy (usukidoll):

take the derivative in terms of x first and then do it again in terms of x for Fxx. treat y as a constant. leave it alone

OpenStudy (jhannybean):

Yeah but isnt the second partial going to = 0?

OpenStudy (usukidoll):

ya

OpenStudy (usukidoll):

Fx = 1 Fxx = 0

OpenStudy (jhannybean):

then multiplying it to fyy would make the distance 0...technically.. fxy=0 and fxxfyy = 0 as well O_o

OpenStudy (usukidoll):

yeahhhhhh... : /

OpenStudy (jhannybean):

my answer is supposed to be (something)/7 xD

OpenStudy (usukidoll):

w ell tell the prof that you tried it and got o's in retgurn

OpenStudy (jhannybean):

Haha thing is she's already solved it...hmm. but wait.-feeling lazy to type it all out-

OpenStudy (jhannybean):

\[\large d= \sqrt{(x-1)^2+(y-2)^2+(z-3)^2}\]\[\ \ \ \large -3z=x+2y-6\]\[\ \ \ \ \ \ \ \large z=-\frac{x}{3} -\frac{2}{3}y+2\]\[\large d=\sqrt{(x-1)^2+(y-2)^2+(-\frac{1}{3}x-\frac{2}{3}y-1)^2}\]\[\large d^2=(x-1)^2 +(y-2)^2 +(-\frac{1}{3}x-\frac{2}{3}y-1)^2\] and now we solve for\( \large f_x \) and \(\large f_y\)?

OpenStudy (usukidoll):

chain rule... and only take derivatives in terms of x

OpenStudy (anonymous):

geomtrical but long method x+2y+3z=6 1,2,3 direction cosines of line normal to this plane which passes through 1,2,3 (x-1)/1=(y-2)/3=(z-3)/3 =>x=y/2=z/3 =k putting in eqn of plane gives point k+4k+9k=14k=6 =>k=6/14=3/7 x=3/7, y=6/7, z=9/7 distance= sqrt((1-3/7)^2+(2-6/7)^2+(3-9/7)^2)=ans

OpenStudy (anonymous):

point is (3/7,6/7,9/7)

OpenStudy (jhannybean):

Could you explain step 2? How did you find "(x-1)/1=(y-2)/3=(z-3)/3 =>x=y/2=z/3 =k putting in eqn of plane gives point k+4k+9k=14k=6 =>k=6/14=3/7"?

OpenStudy (jhannybean):

Oh, so we're trying to use parametric equations to find the point on the tangent line

OpenStudy (anonymous):

yes